a jar of nickels and dimes contains 5.45. there are 8 more dimes than nickles. How many of each are there?
Answer by checkley77(12844)

D=N+8
.05N+.10D=5.45
.05N+.10(N+8)=5.45
.05N+.10N+.8=5.45
.15N=5.45-.8
.15N=4.65
N=4.65/.15
N=31 IS THE NUMBER OF NICKELS.
D=31+8=39 DIMES.
PROOF:
.05*31+.10*39=5.45
1.55+3.90=5.45
5.45=5.45